# Wikontic: Constructing Wikidata-Aligned, Ontology-Aware Knowledge Graphs with Large Language Models

### v1

@@ Abstract
Alla Chepurova1,2   Aydar Bulatov1,2  Yuri Kuratov1,2  Mikhail Burtsev3 1Cognitive AI Systems Lab, Moscow, Russia 2Moscow Independent Research Institute of Artificial Intelligence, Moscow, Russia 3London Institute for Mathematical Sciences, London, UK {chepurova,bulatov,kuratov}@cogailab.com, mb@lims.ac.uk
@@ 1 Introduction
A substantial amount of the world’s knowledge exists solely in unstructured textual form, such as news, scientific articles, blogs and posts on social networks. While large language models (LLMs) are capable of extracting insights from these data, their internal representations are latent and often unverifiable, making them susceptible to hallucinations. In contrast, knowledge graphs (KGs) record information as explicit subject-relation-object triplets, supporting verifiable queries, incremental updates, and multi-step reasoning (e.g., in answering compositional questions), making KGs a reliable complement to both LLMs and retrieval-augmented generation (RAG) systems. Therefore, creating high-quality KGs directly from raw text provides reliable, transparent knowledge that complements LLMs and RAG systems.
@@
Extracting structured knowledge from text is a long-standing challenge in Information Retrieval. A common formulation is closed information extraction (cIE), which assumes a predefined ontology with fixed sets of entity types and relation predicates drawn from an existing KG, and seeks to recover all triplets that conform to that schema. Classical cIE pipelines decompose the task into stages such as named entity recognition and relation classification zeng2014relation; zhang2020minimize. However, separating these stages leads to error accumulation and prevents sharing of information across tasks. More recent end-to-end approaches cast extraction as a sequence-to-sequence problem, training models to directly generate triplets from text distiawan2019neural; cabot2021rebel; josifoski2021genie or complete missing links in KG yao2019kg. While this reduces error propagation, such models remain difficult to adapt to new domains, requiring costly retraining on high-quality annotated corpora that remain scarce. LLMs offer a promising alternative: their broad knowledge and strong prompting abilities enable open-domain extraction without expensive task-specific training wang2020language; josifoski2023exploiting; chepurova2024prompt.
@@
In contrast to cIE, open information extraction (oIE) does not impose predefined entity and relation names, ontology constraints, and constructs KGs from scratch. This flexibility makes oIE an attractive tool for augmenting LLMs and RAG systems, with recent work demonstrating reduced inference costs and more reliable retrieval chen2024sac; gutierrez2024hipporag; guo2024lightrag; li2024graphreader; han2024retrieval; gutierrez2025rag. For instance, AriGraph anokhin2024arigraph learns KGs to create semantic long-term memory, while Distill-SynthKG choubey2024distill integrates mixed text–triplet structures to improve question answering. However, most current oIE pipelines rely on KGs mainly as auxiliary scaffolds for structuring text retrieval, rather than treating the KG itself as a high-quality knowledge resource. This perspective overlooks the potential of compact and non-redundant KGs to represent information directly, and leaves their quality and reasoning capabilities underexploited. As a result, the practical use of oIE KGs remains limited. Extracted triplets often contain heterogeneous surface forms—for instance, “NYC located in USA” versus “New York City in-country United States”—fragmenting the KG into redundant or inconsistent representations. With scale, synonymy, coreference, and predicate variation accumulate, eroding the very strengths that motivate KG construction in the first place: precision, interpretability, and logical consistency.
@@
To address this, we combine the flexibility of oIE with the structural rigor of cIE by leveraging external ontologies. Wikidata vrande2012wikidata, one of the largest community-maintained knowledge bases, offers rich entity classes, relation schemas, and domain–range constraints across more than 100M entities. Its breadth allows coverage from common sense to specialized domains, while its formal constraints provide principled supervision for validating LLM outputs. Yet, integrating such ontology guidance into a fully automated pipeline poses key challenges: (i) extracting candidate triplets without predefined labels, (ii) typing and disambiguating entities under ontology classes despite lexical ambiguity, and (iii) refining nodes and edges iteratively while preserving alignment.
@@
In this paper, we address these challenges with Wikontic, a multi-stage framework that constructs Wikidata-aligned, ontology-aware KGs directly from text using LLMs. Unlike prior work that apply Wikidata ontology only for evaluation or entity linking polat2025testing, we integrate a large-scale ontology from Wikidata directly into the information extraction pipeline. Wikontic includes six components: (i) a curated ontology database derived from Wikidata, (ii) candidate triplet extraction with qualifiers, (iii) ontology-aware triplet refinement enforcing schema constraints, (iv) subject/object name refinement for entity deduplication, (v) KG storage, and (vi) retrieval for multi-hop question answering. Starting from unstructured text, Wikontic extracts triplets with LLMs, types and validates them against Wikidata, and deduplicates entities to yield compact, consistent KGs.
@@
The resulting KGs are both interpretable and effective. When used as the sole knowledge source for multi-hop QA, Wikontic achieves competitive performance with RAG and KG-based methods that rely on raw text as context lee2024human; li2024graphreader; anokhin2024arigraph; panda2024holmes; gutierrez2025rag. Moreover, our graphs exhibit superior coverage of salient information and strong internal connectivity between relevant nodes. In summary, our main contributions are:
@@
1. We introduce Wikontic, which (1) extracts candidate triplets, (2) enforces schema and ontology constraints on types and domain–range, (3) and performs alias-aware entity normalization and deduplication to reduce redundancy, yielding ontology-consistent KGs.
@@
2. We show that Wikontic’s KGs are ontology-consistent and low redundancy with strong coverage and connectivity; on MuSiQue, the correct answer entity is present in 96% of generated triplets.
@@
3. Using the KG as the sole knowledge source (no access to the original text) on multi-hop question answering, Wikontic attains 76.0 F1 on HotpotQA and 59.8 F1 on MuSiQue, matching or surpassing several RAG/KG baselines that still rely on text.
@@
4. Wikontic achieves state-of-the-art results on the MINE-1 benchmark, reaching 86% information-retention scores.
@@ 2 Methods: Wikontic
Figure 1: Overview of Wikontic: an ontology-guided pipeline that constructs a Wikidata-aligned KG from text. (1) An LLM extracts candidate (subject, relation, object) triplets (gray). (2) The extracted triplets are then refined using Wikidata’s ontology: entity types are assigned (colored nodes), and relations that violate ontology constraints are corrected or removed. (3) Finally, entities names are normalized and duplicate surface forms are merged. The resulting graph is de-duplicated, ontology-consistent, and ready for downstream tasks.
@@
To enable triplet validation and alignment, the pipeline stores ontology rules and current KG (Section 2.1). The triplet extraction pipeline (Section 2.2) consists of three main stages: (i) triplet candidate extraction with contextual metadata, (ii) ontology-aware triplet refinement, and (iii) entity normalization and deduplication. These stages aim to enforce structural validity and reduce redundancy, to produce cleaner and more semantically coherent KG that can replace raw text in RAG for multi-hop QA tasks (Section 2.3).
@@ 2.1 Ontology and KG Databases
To support constraint generalization, we recursively expanded entity types using ’instance of’ (P31) and ’subclass of’ (P279) relations, building full taxonomies from each type up to the root. Such hierarchy is essential because relation constraints are defined at different levels of abstraction. For example, a relation may allow connections between instances of the broader class ’audiovisual work’, even if the entity is typed more specifically as ’film’. By propagating the allowed properties of each parent type downwards to its children entity type, we ensure that entities can still be matched to valid relations whenever the constraint applies to its parent class.
We collected labels and aliases for all entity types and relations. Dense retrieval indexes for relation and entity type names, as well as their respective aliases, support semantic search. These indexes allow us to semantically align relation and entity types names from extracted triplets with Wikidata definitions, even when surface forms differ.
@@
Dense retrieval indexes used in both databases were built with Contriever embeddings contriever and Atlas MongoDB vector search [^1] https://www.mongodb.com/products/platform/atlas-vector-search. MongoDB’s hybrid support for structured queries and dense retrieval enables both efficient graph and semantic search.
@@ 2.2 Ontology-aware Triplet Extraction
However, LLM outputs may be semantically redundant or structurally inconsistent. The names of entity and relation may not align with existing entities and relations already present in the KG. For example, an LLM might extract "Nolan" as the subject from one input text and extract "Christopher Nolan" from the other one; or relations like "directed" vs. "director" might appear in different grammatical forms or inverse directions in different input texts (see Figure 1, bottom). Without additional correction, these inconsistencies lead to entity duplication and increased KG size, which degrades both storage efficiency and downstream reasoning. Thus, to improve consistency and reduce redundancy, next steps of the pipeline validate extracted triplets using Wikidata’s ontology and normalize entity names.
@@
Relation validation: Using Wikidata constraints, we identify all relations that can legally connect the subject and object types, including inverse combinations of subject and object types (e.g., directed vs. director). Candidate relations are ranked by cosine similarity to the originally extracted relation.
@@
This step aims to ensure that the resulting KG is compact by avoiding redundant entities with different surface form and evolving by supporting incremental updates with discovery on new entities. A detailed example for the second step is provided on Figure 6 in Appendix A.6.
@@ 2.3 Retrieval for QA
Given a question, the LLM decomposes it into the first 1-hop subquestion. For each subquestion (1) the LLM identifies explicitly mentioned or potentially relevant entities; (2) links extracted entities to KG nodes and selects those most relevant for the current step; (3) given the retrieved subgraph formed by the neighborhood of the selected entities, the LLM generates an answer to the subquestion; (4) conditioned on the previous answer, the LLM formulates the next subquestion. This iterative process continues for up to five subquestions, after which the LLM produces the final answer. Implementation details and prompt templates are provided in Appendix A.5.
@@ 2.4 Evaluation
Existing benchmarks for triplet extraction suffer from substantial limitations: annotated closed IE datasets are small-scale, noisy, and often incomplete josifoski2023exploiting; josifoski2021genie; huguet-cabot-navigli-2021-rebel, while open IE corpora are difficult to align with real-world KGs and provide unreliable ground truth for evaluation stanovsky2018supervised. Constructing high-quality datasets is costly, as annotators must not only identify all explicit and implicit entity and relation mentions but also align them to the complex schemas of large KGs such as Wikidata, which contain thousands of entity types and relations. Recently, the MINE benchmark mo2025kggen was introduced to address some of these issues by evaluating KGs through information retention rather than exact triplet-level supervision. We evaluated Wikontic on the MINE-1 task and observed much higher information retention performance. However, while MINE provides a useful and scalable proxy for KG quality, it does not include complete ground-truth triplets. Therefore, it cannot measure classical precision or recall and instead captures only the degree to which a constructed KG preserves factual information.
To further address this, we adopt an alternative evaluation strategy. We measure KG quality through (1) structural compactness (i.e., non-redundancy and deduplication) and (2) performance in downstream multi-hop QA. In this setup, the LLM must answer factual questions in a text-free setting using only the constructed KG, without access to the original source texts, unlike retrieval-augmented methods such as HippoRAG gutierrez2024hipporag, AriGraph anokhin2024arigraph, and Holmes panda2024holmes. This design makes QA a functional proxy for two key properties of an extracted KG: (a) factual correctness, since noisy or invalid triplets directly impede correct answers, and (b) coverage and completeness, since incomplete graphs restrict multi-hop reasoning. Despite existing advances in aligning KGs with LLMs and adapting for QA han2022self; dai2025large; sui2024can; pan2024unifying, we deliberately refrain from training additional models to estimate the KG quality itself.
@@
We perform evaluations on KGs extracted using different LLMs: gpt-4.1, gpt-4.1-mini, gpt-4o-mini [^2] https://platform.openai.com/docs/models, and Llama-3.3-70b-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, and assess the quality of the resulting KG on two multi-hop QA datasets: MuSiQue trivedi2022musique and HotpotQA yang2018hotpotqa. We used the same questions and candidate passages, including both supporting and distractor passages, used in HippoRAG gutierrez2024hipporag and AriGraph anokhin2024arigraph to compare the results with existing methods.
@@ 3.1 Evaluations on MINE-1
We evaluated Wikontic on the MINE-1 benchmark, which measures how much factual information from the source text is retained in the constructed KGs using an LLM-as-a-judge protocol from the original study mo2025kggen. Figure 2 displays the retention scores distribution in articles of MINE-1 for KGGen, GraphRAG, and Wikontic. Table 1 demonstrates the results for both KGGen and Wikontic with different LLM backbones. Wikontic consistently outperforms KGGen, reaching 84% with gpt-4o and 86% with gpt-4.1-mini, compared to KGGen’s best score of 73% (Claude Sonnet 3.5). These results demonstrate that Wikontic effectively preserves factual information during the construction of the KG.
@@
Table 1: MINE-1 information-retention scores for KGGen and Wikontic. Wikontic achieves the highest retention performance across all evaluated LLMs.

### v2

@@ Abstract
Alla Chepurova1,2   Aydar Bulatov1,2  Mikhail Burtsev3  Yuri Kuratov1,2 1Cognitive AI Systems Lab, Moscow, Russia 2Moscow Independent Research Institute of Artificial Intelligence, Moscow, Russia 3London Institute for Mathematical Sciences, London, UK {chepurova,bulatov,kuratov}@cogailab.com, mb@lims.ac.uk
@@ 1 Introduction
A substantial amount of the world’s knowledge exists solely in unstructured textual form, such as news, scientific articles, blogs, and posts on social networks. While large language models (LLMs) are capable of extracting insights from these data, their internal representations are latent and often unverifiable, making them susceptible to hallucinations. In contrast, knowledge graphs (KGs) record information as explicit subject-relation-object triplets, supporting verifiable queries, incremental updates, and multi-step reasoning (e.g., in answering compositional questions), making KGs a reliable complement to both LLMs and retrieval-augmented generation (RAG) systems. Therefore, creating high-quality KGs directly from raw text provides reliable, transparent knowledge that complements LLMs and RAG systems.
@@
Extracting structured knowledge from text is a long-standing challenge in Information Retrieval. A common formulation is closed information extraction (cIE), which assumes predefined fixed sets of entities and relation predicates drawn from an existing KG and seeks to recover all triplets that conform to that schema. Classical cIE pipelines decompose the task into stages such as named entity recognition and relation classification Zeng et al. (2014); Zhang et al. (2020). However, separating these stages leads to error accumulation and prevents the sharing of information between tasks. More recent end-to-end approaches approach extraction as a sequence-to-sequence problem, training models to directly generate triplets from text Distiawan et al. (2019); Cabot and Navigli (2021); Josifoski et al. (2022) or complete missing links in KG Yao et al. (2019). While this reduces error propagation, such models remain difficult to adapt to new domains, requiring costly retraining on high-quality annotated corpora that remain scarce. LLMs offer a promising alternative: their broad knowledge and strong prompting abilities enable open-domain extraction without expensive task-specific training Wang et al. (2020); Josifoski et al. (2023); Chepurova et al. (2024).
@@
In contrast to cIE, open information extraction (oIE) does not impose predefined entity and relation names, ontology constraints, and constructs KGs from scratch. This flexibility makes oIE an attractive tool for augmenting LLMs and RAG systems, with recent work demonstrating reduced inference costs and more reliable retrieval Chen et al. (2024); Gutierrez et al. (2024); Guo et al. (2024); Li et al. (2024); Han et al. (2024); Gutiérrez et al. (2025). For example, AriGraph Anokhin et al. (2025) learns KGs to create long-term semantic memory, while Distill-SynthKG Choubey et al. (2024) integrates mixed text–triplet structures to improve question answering. However, most current oIE pipelines rely on KGs mainly as auxiliary scaffolds to structure text retrieval, rather than treating the KG itself as a high-quality knowledge resource. This perspective overlooks the potential of compact and non-redundant KGs to represent information directly, and leaves their quality and reasoning capabilities underexploited. As a result, the practical use of oIE KGs remains limited. Extracted triplets often contain heterogeneous surface forms—for example, “NYC located in USA” versus “New York City in-country United States”—fragmenting the KG into redundant or inconsistent representations. Synonymy, coreference, and predicate variation accumulate with scale, eroding the very strengths that motivate KG construction in the first place: precision, interpretability, and logical consistency.
@@
To address this, we combine the flexibility of oIE with the structural rigor of cIE by leveraging external ontologies. Wikidata (Vrandečić, 2012), one of the largest community-maintained knowledge bases, offers rich entity classes, relation schemas, and domain-range constraints across more than 100M entities. Its breadth allows coverage from common sense to specialized domains, while its formal constraints provide principled supervision for validating LLM outputs. Yet, integrating such ontology guidance into a fully automated pipeline poses key challenges: (i) extracting candidate triplets without predefined labels, (ii) typing and disambiguating entities under ontology classes despite lexical ambiguity, and (iii) refining nodes and edges iteratively while preserving alignment.
@@
In this paper, we address these challenges with Wikontic, a multi-stage framework that constructs Wikidata-aligned, ontology-aware KGs directly from texts using LLMs. Unlike prior works that apply Wikidata ontology only for evaluation or entity linking Polat et al. (2025), we integrate a large-scale ontology from Wikidata directly into the information extraction pipeline. Wikontic includes six components: (i) a curated ontology database derived from Wikidata, (ii) candidate triplet extraction with qualifiers, (iii) ontology-aware triplet refinement enforcing schema constraints, (iv) subject/object name refinement for entity deduplication, (v) KG storage, and (vi) retrieval for multi-hop question answering. Starting from unstructured text, Wikontic extracts triplets with LLMs, types, and validates them against Wikidata, and deduplicates entities to yield compact, consistent KGs.
@@
The resulting KGs are both interpretable and effective. When used as the sole knowledge source for multi-hop QA, Wikontic achieves competitive performance with RAG and KG-based methods that rely on raw text as context Lee et al. (2024); Li et al. (2024); Anokhin et al. (2025); Panda et al. (2024); Gutiérrez et al. (2025). Moreover, our graphs exhibit superior coverage of salient information and strong internal connectivity between relevant nodes. In summary, our main contributions are:
@@
1. We introduce Wikontic, which (1) extracts candidate triplets, (2) enforces schema and ontology constraints on entity types and domain–range, and (3) performs alias-aware entity normalization and deduplication to reduce redundancy, yielding ontology-consistent KGs.
@@
2. We show that Wikontic’s KGs are ontology-consistent, have low redundancy, strong coverage, and connectivity; on MuSiQue, the correct answer entity is present in 96% of generated triplets.
@@
3. Using a KG as the sole knowledge source (no access to the original text) on multi-hop question answering, Wikontic attains 76.0 F1 on HotpotQA and 59.8 F1 on MuSiQue, matching or surpassing several RAG/KG baselines that still rely on text.
@@
4. Wikontic achieves state-of-the-art results on the MINE-1 benchmark, reaching 86% information-retention score.
@@ 2 Methods: Wikontic
Figure 1: Overview of Wikontic: an ontology-guided pipeline that constructs a Wikidata-aligned KG from text. (1) An LLM extracts candidate (subject, relation, object) triplets (gray). (2) The extracted triplets are then refined using Wikidata’s ontology: entity types are assigned (colored nodes), and relations that violate ontology constraints are corrected or removed. (3) Finally, entity names are normalized, and duplicated surface forms are merged. The resulting graph is de-duplicated, ontology-consistent, and ready for downstream tasks.
@@
To enable triplet validation and alignment, the pipeline stores ontology rules and the current KG (Section 2.1). The triplet extraction pipeline (Section 2.2) consists of three main stages: (i) triplet candidate extraction with contextual metadata, (ii) ontology-aware triplet refinement, and (iii) entity normalization and deduplication. These stages aim to enforce structural validity and reduce redundancy, to produce a cleaner and more semantically coherent KG that can replace raw text in RAG for multi-hop QA tasks (Section 2.3).
@@ 2.1 Ontology and KG Databases
To support constraint generalization, we recursively expanded entity types using ’instance of’ (P31) and ’subclass of’ (P279) relations, building full taxonomies from each type up to the root. Such a hierarchy is essential because relation constraints are defined at different levels of abstraction. For example, a relation may allow connections between instances of the broader class ’audiovisual work’, even if the entity is typed more specifically as ’film’. By propagating the allowed properties of each parent type downwards to its children entity type, we ensure that entities can still be matched to valid relations whenever the constraint applies to its parent class.
We collected labels and aliases for all entity types and relations. Dense retrieval indexes for relation and entity type names, as well as their respective aliases, support semantic search. These indexes allow us to semantically align relation and entity type names from extracted triplets with Wikidata definitions, even when surface forms differ.
@@
Dense retrieval indexes used in both databases were built with Contriever embeddings Izacard et al. (2022) and Atlas MongoDB vector search [^1] https://www.mongodb.com/products/platform/atlas-vector-search. MongoDB’s hybrid support for structured queries and dense retrieval enables both efficient graph and semantic search.
@@ 2.2 Ontology-aware Triplet Extraction
However, LLM outputs may be semantically redundant or structurally inconsistent. The entity and relation names may not align with existing entities and relations already present in the KG. For example, an LLM might extract "Nolan" as the subject from one input text and extract "Christopher Nolan" from the other one; or relations like "directed" vs. "director" might appear in different grammatical forms or inverse directions in different input texts (see Figure 1, bottom). Without additional correction, these inconsistencies lead to entity duplication and increased KG size, which degrades both storage efficiency and downstream reasoning. Thus, to improve consistency and reduce redundancy, the next steps of the pipeline validate extracted triplets using Wikidata’s ontology and normalize entity names.
@@
Relation validation: Using Wikidata constraints, we identify all relations that can legally connect selected subject and object types, including inverse combinations of subject and object types (e.g., directed vs. director). These candidate relations are ranked by cosine similarity to the originally extracted relation.
@@
This step aims to ensure that the resulting KG is compact by avoiding redundant entities with different surface forms and evolving by supporting incremental updates with the discovery of new entities. A detailed example for the second step is provided in Figure 6 in Appendix A.6.
@@ 2.3 Retrieval for QA
Given a question, the LLM decomposes it into the first 1-hop subquestion. For each subquestion the LLM (1) identifies explicitly mentioned or potentially relevant entities; (2) links extracted entities to KG nodes and selects those most relevant for the current step; (3) given the retrieved subgraph formed by the neighborhood of the selected entities, generates an answer to the subquestion; (4) conditioned on the previous answer, the LLM formulates the next subquestion. This iterative process continues for up to five subquestions, after which the LLM produces the final answer. Implementation details and prompt templates are provided in Appendix A.5.
@@ 2.4 Evaluation
Existing benchmarks for triplet extraction suffer from substantial limitations: annotated closed IE datasets are small-scale, noisy, and often incomplete Josifoski et al. (2023, 2022); Huguet Cabot and Navigli (2021), while open IE corpora are difficult to align with real-world KGs and provide unreliable ground truth for evaluation Stanovsky et al. (2018). Constructing high-quality datasets is costly, as annotators must not only identify all explicit and implicit entity and relation mentions but also align them to the complex schemas of large KGs such as Wikidata, which contain thousands of entity types and relations. Recently, the MINE benchmark Mo et al. (2025) was introduced to address some of these issues by evaluating KGs through information retention rather than exact triplet-level supervision. We evaluated Wikontic on the MINE-1 task and observed much higher information retention performance. However, while MINE provides a useful and scalable proxy for KG quality, it does not include complete ground-truth triplets. Therefore, it cannot measure classical precision or recall and instead captures only the degree to which a constructed KG preserves factual information.
To further address this, we adopt an alternative evaluation strategy. We measure KG quality through (1) structural compactness (i.e., non-redundancy and deduplication) and (2) performance in downstream multi-hop QA. In this setup, the LLM must answer factual questions in a text-free setting using only the constructed KG, without access to the original source texts, unlike retrieval-augmented methods such as HippoRAG Gutierrez et al. (2024), AriGraph Anokhin et al. (2025), and Holmes Panda et al. (2024). This design makes QA a functional proxy for two key properties of an extracted KG: (a) factual correctness, since noisy or invalid triplets directly impede correct answers, and (b) coverage and completeness, since incomplete graphs restrict multi-hop reasoning. Despite existing advances in aligning KGs with LLMs and adapting for QA Han and Shareghi (2022); Dai et al. (2025); Sui et al. (2025); Pan et al. (2024), we deliberately refrain from training additional models to estimate the KG quality itself.
@@
We perform evaluations on KGs extracted using different LLMs: gpt-4.1, gpt-4.1-mini, gpt-4o-mini [^2] https://platform.openai.com/docs/models, and Llama-3.3-70b-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, and assess the quality of the resulting KG on two multi-hop QA datasets: MuSiQue Trivedi et al. (2022) and HotpotQA Yang et al. (2018). We used the same questions and candidate passages, including both supporting and distractor passages, used in HippoRAG Gutierrez et al. (2024) and AriGraph Anokhin et al. (2025) to compare the results with existing methods.
@@ 3.1 Evaluations on MINE-1
We evaluated Wikontic on the MINE-1 benchmark, which measures how much factual information from the source text is retained in the constructed KGs using an LLM-as-a-judge protocol from the original study Mo et al. (2025). Figure 2 displays the retention scores distribution in articles of MINE-1 for KGGen, GraphRAG, and Wikontic. Table 1 demonstrates the results for both KGGen and Wikontic with different LLM backbones. Wikontic consistently outperforms KGGen, reaching 84% with gpt-4o and 86% with gpt-4.1-mini, compared to KGGen’s best score of 73% (Claude Sonnet 3.5). These results demonstrate that Wikontic effectively preserves factual information during the construction of the KG.
@@
Table 1: MINE-1 information-retention scores for KGGen, GraphRAG, and Wikontic. Wikontic achieves the highest retention performance across all evaluated LLMs.